# Towards a clinically accessible radiology foundation model: open-access and lightweight, with automated evaluation

### v4

@@ Abstract
The scaling laws and extraordinary performance of large foundation models motivate the development and utilization of such models in biomedicine. However, despite early promising results on some biomedical benchmarks, there are still major challenges that need to be addressed before these models can be used in real-world clinics. Frontier general-domain models such as GPT-4V still have significant performance gaps in multimodal biomedical applications. More importantly, less-acknowledged pragmatic issues, including accessibility, model cost, and tedious manual evaluation make it hard for clinicians to use state-of-the-art large models directly on private patient data. Here, we explore training open-source small multimodal models (SMMs) to bridge competency gaps for unmet clinical needs in radiology. To maximize data efficiency, we adopt a modular approach by incorporating state-of-the-art pre-trained models for image and text modalities, and focusing on training a lightweight adapter to ground each modality to the text embedding space, as exemplified by LLaVA-Med. For training, we assemble a large dataset of over 697 thousand radiology image-text pairs. For evaluation, we propose CheXprompt, a GPT-4-based metric for factuality evaluation, and demonstrate its parity with expert evaluation. For best practice, we conduct a systematic ablation study on various choices in data engineering and multimodal training. The resulting LLaVA-Rad (7B) model attains state-of-the-art results on standard radiology tasks such as report generation and cross-modal retrieval, even outperforming much larger models such as GPT-4V and Med-PaLM M (84B). Moreover, LLaVA-Rad requires only one day to be trained on over 697 thousand image-text pairs using a standard 8-A100 GPU cluster, allowing further fine-tuning by clinicians using their own data. The inference of LLaVA-Rad is fast and can be performed on a single V100 GPU in private settings, offering a promising state-of-the-art tool for real-world clinical applications.
@@ Introduction
Foundation models, trained on massive amounts of unlabelled data using self-supervised learning, enable rapid adaptation to various downstream tasks with minimal requirement for task-specific labeled data [5, 36]. Due to the high cost of annotating biomedical data [22, 21], foundation models are poised to become a new paradigm in biomedicine, achieving state-of-the-art results on many applications, including medical question answering [47, 36] and medical image classification [4, 3]. Recently, multimodal generative AI has emerged as an exciting frontier in the biomedical domain, expanding the application scope from single-modality to multi-modality (e.g., text and image), such as visual question answering and radiology report generation [29, 47, 44]. While existing models are still largely evaluated on artificial biomedical benchmarks, their promising performance demonstrates their potential in clinical applications.
@@
However, there are still major bottlenecks hindering the progress toward using foundation models in real-world clinical settings. First, sharing patient data with large foundation models hosted on the cloud is subject to privacy concerns [45]. Therefore, clinicians may prefer to run inference and fine-tuning locally. Second, existing state-of-the-art models are often very large and resource-intensive, which makes local deployment challenging. Smaller models incur smaller carbon footprint [46] and offer reduced serving costs and latency, which is of particular importance in resource-constrained settings outside of data centers [51]. However, while small language models have shown success in text domains [16, 34, 28], small multimodal models still have significant performance gaps compared to larger models [47, 44]. Third, many state-of-the-art models are inaccessible [51], necessitating the development of effective open-source models for biomedicine. Finally, even the best models are still subject to errors such as hallucination and there exist limited methods to reliably verify the correctness of model outputs at scale, especially in the specialized field of biomedicine [10].
@@
To bridge this gap between existing medical foundation models and real-world clinical applications, we have developed LLaVA-Rad, a small multimodal model (SMM) that attains state-of-the-art performance in standard radiology imaging tasks (Fig. 1). We focus our study on identifying key findings from chest X-ray (CXR) images, the most commonly performed medical imaging examination. Automatically drafting high-quality radiology reports is a challenging but clinically meaningful task that could directly increase radiologist productivity and potentially improve communication and decrease burnout [27]. Existing frontier models such as GPT-4V still have a large performance gap even on such a fundamental medical application. To develop LLaVA-Rad, we adopt a modular approach by incorporating state-of-the-art open-source pretrained models for image and text modalities, and focusing on training a lightweight adapter to ground each modality to the text embedding space, as exemplified by LLaVA-Med [29].
@@ Results
We first evaluated LLaVA-Rad on the widely-used radiology report generation benchmark MIMIC-CXR (Fig.  2, Supplementary Figure  1) using metrics assessing lexical similarity and factual accuracy. In particular, lexical similarity metrics, such as BLEU and ROUGE, are used in traditional NLP to assess the model’s ability to produce contextually and stylistically aligned output. On the other hand, factual accuracy metrics, including CheXbert-based and RadGraph-based F1 scores [43, 25] are more clinically relevant because they gauge the extent to which the generated reports accurately reflect imaging findings. A high-quality radiology report necessitates both coherence and factual accuracy, underscoring the importance of evaluating these two types of metrics to ensure the report is both intelligible and clinically precise. Any inaccuracies in the report can potentially lead to adverse patient outcomes, emphasizing the critical need for factual correctness.
We found that LLaVA-Rad achieved superior performance on both groups of metrics (Fig. 2A-D). When compared to other models of equivalent size (7B parameters), such as LLaVA-Med [29], CheXagent [8] and MAIRA-1 [23], LLaVA-Rad demonstrates significant advancements in performance across all evaluated metrics. Furthermore, LLaVA-Rad is more efficient than the current overall leading model, Med-PaLM M [47], with an order of magnitude fewer parameters. This efficiency does not come at the cost of effectiveness; LLaVA-Rad outperforms Med-PaLM M in the most important existing lexical similarity and factual correctness metrics for radiology text (ROUGE-L and F1-RadGraph [54]). A more detailed evaluation (Supplementary Table 1) shows that Med-PaLM M marginally surpasses LLaVA-Rad by F1-5 CheXbert metrics, which assess only a small subset of 5 potential abnormalities, and the performance gap is minimal. Most of these competing models also use MIMIC-CXR for training (with the notable exception of LLaVA-Med). We attribute the promising performance of LLaVA-Rad to its modular design, which is more data efficient. The efficiency and the high degree of factual and lexical precision of LLaVA-Rad demonstrate its potential in real-world applications where large models are computationally too costly.
@@
It is well known that existing ngram-based automated report evaluation methods might be biased to pre-defined conditions and have limited correlation with expert assessments [52]. We thus explore the utility of an LLM-based evaluation system, which has shown success in other domains [32, 48, 15]. Specifically, we employ GPT-4 as an evaluator to count how often the generated report commits errors in each of the following six categories, as per a previous study [52]: false positive finding, omission of finding, incorrect location/position of finding, incorrect severity of finding, mention of comparison that is not present in the reference report, omission of comparison describing a change from a previous study. For each error type, we further instruct GPT-4 to distinguish clinically significant and clinically insignificant errors.
@@
We also perform a head-to-head comparison of the calculation of total errors as determined by GPT-4Turbo with manual expert ratings in a leave-one-rater-out fashion. Fig. 4A summarizes the results of this comparison, which on average shows a mean absolute difference (MAD) of 0.71 between the left-out rater and the average of the remaining ones, whereas GPT-4T has on average 0.55 MAD. We find that the MAD between GPT-4T CheXprompt score and the left-in expert average is smaller compared to the left-out expert in 3 out of 6 cases (P<0.001), slightly larger in 2 out of 6 cases (P<0.001), and not significantly different (P>0.05) in the remaining one case. Altogether, we find that GPT-4T is indistinguishable from expert raters in calculating the total number of errors, increasing our confidence in using this proposed automated metric as a new evaluation method that directly aligns with expert opinions.
After assuring the effectiveness of the GPT-based metric, we evaluated the performance of LLaVA-Rad on the held-out MIMIC-CXR test set using CheXprompt (Fig. 4C,D). Similar to our observation using existing metrics, LLaVA-Rad outperforms publicly available report generation models, generating fewer clinically significant and total errors compared to GPT-4V and CheXagent. Moreover, by comparing models within the LLaVA family (e.g., LLaVA-Rad, LLaVA-Med, LLaVA), we observed that fewer errors are made in the generated reports as increasing domain specialization is performed. In particular, LLaVA-Rad generates fewer errors than LLaVA-Med, a LLaVA model tailored to medicine, and LLaVA-Med generates fewer errors than the general-domain model LLaVA. This suggests a trade-off between domain-specific performance and broad applicability, supporting our intuition of developing LLaVA-Rad by continue-pretraining a general model using large amounts of domain-specific data. Finally, to determine the clinical utility of LLaVA-Rad, we explore using the percentage of error-free reports to track the overall performance of report-generation models. A higher percentage of error-free reports increases the utility of a report generation model, given that it directly reflects the number of reports that require little to no radiologist modification following automated generation. Notably, LLaVA-Rad has the highest percentage of error-free reports, with 6.79% reports free of clinically significant errors, and 2.58% free of errors. The same trend was observed in the external validation datasets (CheXpert and Open-I), where LLaVA-Rad demonstrated fewer clinically significant and total errors (Fig. 3), with up to 26% error-free reports. While there remains undoubtedly significant growth space, this direction and latest progress are promising.
@@
Conducting thorough ablation studies for LLMs is often intractable due to the costly training of multiple variants. In contrast, the small size of LLaVA-Rad enables us to efficiently conduct ablation studies that explain the promising performance of LLaVA-Rad and potentially help derisk design choices for larger models. We compared LLaVA-Rad with 8 variants described in Supplementary Table 2. In Fig.  5A,B, we investigate two key technical ideas used in LLaVA-Rad: the effect of pre-training a domain-specific image encoder using 697 thousand diverse CXR image-text pairs and the effect of using GPT-4 to augment and organize the data. First, to understand the effect of pre-training an image encoder, we compare LLaVA-Rad with three increasingly domain-specific variants: an image encoder from OpenAI CLIP, an image encoder using BiomedCLIP, and an image encoder from BiomedCLIP but continually pre-trained using MIMIC-CXR only. To avoid data contamination, we only used the training split of MIMIC-CXR. We did not find noticeable overlap between MIMIC-CXR training split and the PubMed data used to pre-train BiomedCLIP. We found that the MIMIC-CXR-based image encoder outperforms the other two variants, indicating the effectiveness of training a domain-specific image encoder. BiomedCLIP-CXR outperforms the BiomedCLIP image encoder continue-pretrained only on MIMIC-CXR (BiomedCLIP-MIMIC-CXR), illustrating the advantage in pre-training using more diverse CXR datasets. Second, we studied the effect of using GPT-4 to process and augment the MIMIC-CXR report data. Supplementary Table 3 summarizes the data that LLaVA-Rad uses for training in the second and the third stages. It is a combination of rule-based and GPT-structured data. We compare LLaVA-Rad with a variant that only uses rule-based data and a variant that only uses GPT-structured data. We found that LLaVA-Rad attains a better performance than both variants, indicating the effectiveness of GPT-4 data augmentation. The variant that only uses GPT-structured data outperforms the one that only uses rule-based data on factual accuracy metrics, confirming the effectiveness of GPT-4-based structuring in generating clinically precise reports. Finally, it is expected that rule-based variant outperforms GPT-structured variant on n-gram lexical metrics, because the test data is also from rule-based data. These ablation studies support our intuition that domain-specific data can help us build a small but effective domain-specific model, and help inform best practice in training larger models.
@@ Discussion
To address the significant challenges of developing foundation models for real-world clinical settings, our work introduces LLaVA-Rad, a lightweight radiology small multimodal model (SMM) that offers open-source accessibility while attaining new state-of-the-art results in the domain of radiology report generation. By curating a dataset of 697 thousand CXR images paired with radiology reports from diverse sources, using GPT-4 for dataset processing, coupled with a modular three-stage curriculum training method, we have developed a model that outperforms its larger counterparts, such as GPT-4V and Med-PaLM M, and demonstrates exceptional proficiency in generating accurate and lexically similar radiology reports on the evaluation datasets. Through our attention visualization techniques, LLaVA-Rad offers deep insights into how it prioritizes key regions in chest X-rays, correlating them with specific findings in the generated reports. Furthermore, our utilization of CheXprompt for evaluation demonstrates a closer alignment of automated scoring with expert radiologist assessments than traditional metrics, reinforcing LLaVA-Rad’s superiority in clinical report generation.
@@
While LLaVA-Rad represents a significant advancement in medical foundation models, our research acknowledges several areas for future exploration and improvement. First, the current scope of LLaVA-Rad is limited to CXRs. While CXR is the most common medical image examination, future iterations should evaluate the feasibility of our method on alternative anatomies (e.g., abdomen or extremities) and modalities (e.g., computed tomography or ultrasound) to enhance the model’s applicability and utility across diverse application scenarios. Attention-based attribution methods have been found to be more effective at explaining model decisions and to be more useful by radiologists [50], and our attention visualization technique does appear to highlight sensible patterns. However, alternative saliency-based methods for CXR interpretation algorithms such as Grad-CAM have been shown to have limited correlation with expert assessments and limited robustness to input perturbations [42, 55]. There is a pressing need for a more exhaustive evaluation of such grounding strategies. These would further improve the model’s explainability and interpretability, making it more transparent and trustworthy for clinical use. Another consideration is the inherently multimodal nature of modern medical practice, which integrates various patient information streams, including historical medical images, medical records, lab tests, and vital signs. Integrating these diverse and longitudinal data sources into medical foundation models like LLaVA-Rad could significantly enrich the model’s understanding and analysis, leading to more nuanced and holistic patient assessments.
LLaVA-Rad exemplifies a significant leap toward making advanced clinical reasoning capabilities accessible with limited computational resources, thus paving the way for broader clinical applications and impact. The pursuit of open-source, lightweight, high-performing models that not only extends to various medical imaging types but also incorporates multimodality and interpretability, embodies the next frontier in medical foundation model development. Such advancements will bridge the gap between current technological capabilities and the real-world demands in clinical applications, moving us closer to achieving meaningful improvements in patient outcomes.

### v5

# Towards a clinically accessible radiology multimodal model: open-access and lightweight, with automatic evaluation

@@ Abstract
The scaling laws and extraordinary performance of large foundation models motivate the development and utilization of such models in biomedicine. However, despite early promising results on some biomedical benchmarks, there are still major challenges that need to be addressed before these models can be used in real-world clinics. Frontier general-domain models such as GPT-4V still have significant performance gaps in multimodal biomedical applications. More importantly, less-acknowledged pragmatic issues, including accessibility, model cost, and tedious manual evaluation make it hard for clinicians to use state-of-the-art large models directly on private patient data. Here, we explore training open-source small multimodal models (SMMs) to bridge competency gaps for unmet clinical needs in radiology. To maximize data efficiency, we adopt a modular approach by incorporating state-of-the-art pre-trained models for image and text modalities, and focusing on training a lightweight adapter to ground each modality to the text embedding space. For training, we assemble a large dataset of over 697 thousand radiology image-text pairs. For evaluation, we propose CheXprompt, a GPT-4-based metric for assessing factual accuracy, and demonstrate its parity with expert evaluation. For best practice, we conduct a systematic ablation study on various choices in data engineering and multimodal training. The resulting LLaVA-Rad (7B) model attains state-of-the-art results on standard radiology tasks such as report generation and cross-modal retrieval, even outperforming much larger models such as GPT-4V and Med-PaLM M (84B). Moreover, LLaVA-Rad requires only one day to be trained on over 697 thousand image-text pairs using a standard 8-A100 GPU cluster, allowing further fine-tuning by clinicians using their own data. The inference of LLaVA-Rad is fast and can be performed on a single V100 GPU in private settings, offering a promising state-of-the-art tool for real-world clinical applications.
@@ Introduction
Foundation models, trained on massive amounts of unlabelled data using self-supervised learning, enable rapid adaptation to various downstream tasks with minimal requirement for task-specific labeled data [5, 36, 52]. Due to the high cost of annotating biomedical data [22, 21], foundation models are poised to become a new paradigm in biomedicine, achieving state-of-the-art results on many applications, including medical question answering [47, 36] and medical image classification [4, 3]. Recently, multimodal generative AI has emerged as an exciting frontier in the biomedical domain, expanding the application scope from single-modality to multi-modality (e.g., text and image), such as visual question answering and radiology report generation [29, 47, 44]. While existing models are still largely evaluated on artificial biomedical benchmarks, their promising performance demonstrates their potential in clinical applications.
@@
However, there are still major bottlenecks hindering the use of foundation models in real-world clinical settings. First, sharing patient data with large foundation models hosted on the cloud is subject to privacy concerns [45]. Therefore, clinicians may prefer to run inference and fine-tuning locally. Second, existing state-of-the-art models are often very large and resource-intensive, which makes local deployment challenging. Smaller models incur smaller carbon footprint [46] and offer reduced serving costs and latency, which is of particular importance in resource-constrained settings outside of data centers [51]. However, while small language models have shown success in text domains [16, 34, 28], small multimodal models (SMMs) still have significant performance gaps compared to larger models [47, 44]. Third, many state-of-the-art models are inaccessible [51], necessitating the development of effective open-source models for biomedicine. Finally, existing evaluation methods of factual correctness exhibit limited correlation with expert assessments [53]. Since even the best models are still subject to errors such as hallucination, it is crucial to develop reliable methods to evaluate the correctness of model outputs at scale, especially in the specialized field of biomedicine [10].
@@
To bridge this gap between existing medical foundation models and real-world clinical applications, we have developed LLaVA-Rad, a SMM that attains state-of-the-art performance in standard radiology imaging tasks (Fig. 1), in addition to CheXprompt, an automatic scoring metric for factual correctness. We focus our study on identifying key findings from chest X-ray (CXR) images, the most commonly performed medical imaging examination. Automatically drafting high-quality radiology reports is a challenging but clinically meaningful task that could directly increase radiologist productivity and potentially improve communication and decrease burnout [27]. Existing frontier models such as GPT-4V still have a large performance gap even on such a fundamental medical application. To develop LLaVA-Rad, we adopt a modular approach by incorporating state-of-the-art open-source pretrained models for image and text modalities, and focusing on training a lightweight adapter to ground each modality to the text embedding space.
@@ Results
We first evaluated LLaVA-Rad on the widely-used radiology report generation benchmark MIMIC-CXR (Fig.  2, Supplementary Figure  1) using metrics assessing lexical similarity and factual accuracy. In particular, lexical similarity metrics, such as BLEU and ROUGE, are used in traditional NLP to assess the model’s ability to produce contextually and stylistically aligned output. On the other hand, factual accuracy metrics, including CheXbert-based and RadGraph-based F1 scores [43, 25] are more clinically relevant because they gauge the extent to which the generated reports accurately reflect imaging findings.
We found that LLaVA-Rad achieved superior performance on both groups of metrics (Fig. 2A-D). When compared to other models of equivalent size (7B parameters), such as LLaVA-Med [29], CheXagent [8] and MAIRA-1 [23], LLaVA-Rad demonstrates significant advancements in performance across all evaluated metrics. Furthermore, LLaVA-Rad is more efficient than the current overall leading model, Med-PaLM M [47], with an order of magnitude fewer parameters. This efficiency does not come at the cost of effectiveness; LLaVA-Rad outperforms Med-PaLM M in the most important existing lexical similarity and factual correctness metrics for radiology text (ROUGE-L and F1-RadGraph [55], with a relative improvement of 12.1% and 10.1% respectively). A more detailed evaluation (Supplementary Table 1) shows that Med-PaLM M marginally surpasses LLaVA-Rad by F1-5 CheXbert metrics, which assess only a small subset of 5 potential abnormalities, and the performance gap is minimal (<1% relative improvement). Most of these competing models also use MIMIC-CXR for training (with the notable exception of LLaVA-Med). We attribute the promising performance of LLaVA-Rad to its modular design, which is more data efficient. The efficiency and the high degree of factual and lexical precision of LLaVA-Rad demonstrate its potential in real-world applications where large models are computationally too costly.
@@
It is well known that existing ngram-based automated report evaluation methods might be biased to pre-defined conditions and have limited correlation with expert assessments [53]. We thus explore the utility of an LLM-based evaluation system, which has shown success in other domains [32, 48, 15]. Specifically, we employ GPT-4 as an evaluator to count how often the generated report contains errors in each of the following six categories, as per a previous study [53]: false positive finding, omission of finding, incorrect location/position of finding, incorrect severity of finding, mention of comparison that is not present in the reference report, omission of comparison describing a change from a previous study. For each error type, we further instruct GPT-4 to distinguish clinically significant and clinically insignificant errors.
@@
We also perform a head-to-head comparison of the calculation of total errors as determined by GPT-4 Turbo with manual expert ratings in a leave-one-rater-out fashion. Fig. 4A summarizes the results of this comparison, which on average shows a mean absolute difference (MAD) of 0.71 between the left-out rater and the average of the remaining ones, whereas GPT-4 Turbo has on average 0.55 MAD. We find that the MAD between GPT-4 Turbo CheXprompt score and the left-in expert average is smaller compared to the left-out expert in 3 out of 6 cases (P<0.001), and not significantly different (P>0.05) in the remaining 3 out of 6 cases. Altogether, we find that GPT-4 Turbo is indistinguishable from expert raters in calculating the total number of errors, increasing our confidence in using this proposed automated metric as a new evaluation method that directly aligns with expert opinions.
After assuring the effectiveness of the GPT-based metric, we evaluated the performance of LLaVA-Rad on the held-out MIMIC-CXR test set using CheXprompt (Fig. 4C,D). Similar to our observation using existing metrics, LLaVA-Rad outperforms publicly available report generation models, generating fewer clinically significant and total errors compared to GPT-4V and CheXagent. Moreover, by comparing models within the LLaVA family (e.g., LLaVA-Rad, LLaVA-Med, LLaVA), we observed that fewer errors are made in the generated reports as increasing domain specialization is performed. In particular, LLaVA-Rad generates fewer errors than LLaVA-Med, a LLaVA model tailored to medicine, and LLaVA-Med generates fewer errors than the general-domain model LLaVA. This suggests a trade-off between domain-specific performance and broad applicability, supporting our intuition of developing LLaVA-Rad by continual-pretraining of a general model using large amounts of domain-specific data. Finally, to determine the clinical utility of LLaVA-Rad, we explore using the percentage of error-free reports to track the overall performance of report-generation models. A higher percentage of error-free reports increases the utility of a report generation model, given that it directly reflects the number of reports that require little to no radiologist modification following automated generation. Notably, LLaVA-Rad has the highest percentage of error-free reports, with 6.79% reports free of clinically significant errors, and 2.58% free of errors. The same trend of improved performance of LLaVA-Rad was observed in the external validation datasets (CheXpert and Open-I), where LLaVA-Rad demonstrated fewer clinically significant and total errors (Fig. 3B,D), with up to 26% error-free reports. CheXprompt illustrates that while there undoubtedly remains room for improvement in automated radiology report generation, the improvement demonstrated by our model is promising.
@@
Conducting thorough ablation studies for LLMs is often intractable due to the costly training of multiple variants. In contrast, the small size of LLaVA-Rad enables us to efficiently conduct ablation studies that explain the promising performance of LLaVA-Rad and potentially inform design choices for larger models. We compared LLaVA-Rad with 8 variants described in Supplementary Table 2. In particular, we investigate two key technical ideas used in LLaVA-Rad: the effect of pre-training a domain-specific image encoder using 697 thousand diverse CXR image-text pairs (Fig.  5A) and the effect of using GPT-4 to augment and organize the data (Fig.  5B). First, to understand the effect of pre-training an image encoder, we compare LLaVA-Rad with three increasingly domain-specific variants: an image encoder from OpenAI CLIP, an image encoder using BiomedCLIP, and an image encoder from BiomedCLIP but continually pre-trained using MIMIC-CXR only (Fig.  5A). To avoid data contamination, we only used the training split of MIMIC-CXR. We did not find noticeable overlap between MIMIC-CXR training split and the PubMed data used to pre-train BiomedCLIP. We found that the MIMIC-CXR-based image encoder outperforms the other two variants, indicating the effectiveness of training a domain-specific image encoder. BiomedCLIP-CXR outperforms the BiomedCLIP image encoder continue-pretrained only on MIMIC-CXR (BiomedCLIP-MIMIC-CXR), illustrating the advantage in pre-training using more diverse CXR datasets. Second, we studied the effect of using GPT-4 to process and augment the MIMIC-CXR report data (Fig.  5B). Supplementary Table 3 summarizes the data that LLaVA-Rad uses for training in the second and the third stages. It is a combination of rule-based and GPT-structured data. We compare LLaVA-Rad with a variant that only uses rule-based data and a variant that only uses GPT-structured data. We found that LLaVA-Rad attains a better performance than both variants, indicating the effectiveness of GPT-4 data augmentation. The variant that only uses GPT-structured data outperforms the one that only uses rule-based data on factual accuracy metrics, confirming the effectiveness of GPT-4-based structuring in generating clinically precise reports. Finally, it is expected that rule-based variant outperforms GPT-structured variant on n-gram lexical metrics, because the test data is also from rule-based data. These ablation studies support our intuition that domain-specific data can help us build a small but effective domain-specific model, and help inform best practice in training larger models.
@@ Discussion
To address the significant challenges of developing foundation models for real-world clinical settings, our work introduces LLaVA-Rad, a lightweight radiology SMM that offers open-source accessibility while attaining new state-of-the-art results in the domain of radiology report generation. By curating a dataset of 697 thousand CXR images paired with radiology reports from diverse sources, using GPT-4 for dataset processing, coupled with a modular three-stage curriculum training method, we have developed a model that outperforms its larger counterparts, such as GPT-4V and Med-PaLM M, and demonstrates exceptional proficiency in generating accurate and lexically similar radiology reports on the evaluation datasets. Through our attention visualization techniques, LLaVA-Rad offers deep insights into how it prioritizes key regions in chest X-rays, correlating them with specific findings in the generated reports. Furthermore, our work introduces CheXprompt, which successfully resolves a major bottleneck of automatic evaluation of factual accuracy of generated radiology reports, by not only demonstrating a closer alignment of automated scoring compared to traditional metrics, but also exhibiting performance on par with expert radiologist annotators. This improved evaluation further illustrates LLaVA-Rad’s superiority in clinical report generation.
@@
While LLaVA-Rad represents a substantial advancement in radiology multimodal models, our research acknowledges several areas for future exploration and improvement. First, the current scope of LLaVA-Rad is limited to CXRs. While CXR is the most common medical image examination, future iterations should evaluate the feasibility of our method on alternative anatomies (e.g., abdomen or extremities) and modalities (e.g., computed tomography or ultrasound) to enhance the model’s applicability and utility across diverse application scenarios. Attention-based attribution methods have been found to be more effective at explaining model decisions and to be more useful by radiologists [50], and our attention visualization technique does appear to highlight sensible patterns. However, alternative saliency-based methods for CXR interpretation algorithms such as Grad-CAM have been shown to have limited correlation with expert assessments and limited robustness to input perturbations [42, 56]. There is a pressing need for a more exhaustive evaluation of such grounding strategies. These would further improve the model’s explainability and interpretability, making it more transparent and trustworthy for clinical use. Another consideration is the inherently multimodal nature of modern medical practice, which integrates various patient information streams, including historical medical images, medical records, lab tests, and vital signs. Integrating these diverse and longitudinal data sources into medical multimodal models like LLaVA-Rad could significantly enrich the model’s understanding and analysis, leading to more nuanced and holistic patient assessments.
LLaVA-Rad exemplifies a significant leap toward making advanced diagnostic capabilities accessible with limited computational resources, thus paving the way for broader clinical applications and impact. The pursuit of open-source, lightweight, high-performing models that not only extends to various medical imaging types but also incorporates multimodality and interpretability, embodies the next frontier in medical multimodal model development. Such advancements will bridge the gap between current technological capabilities and the real-world demands in clinical applications, moving us closer to achieving meaningful improvements in patient outcomes.